# Chaos-based reinforcement learning with TD3

### v1

@@ Abstract
Chaos-based reinforcement learning (CBRL) is a method in which the agent’s internal chaotic dynamics drives exploration. This approach offers a model for considering how the biological brain can create variability in its behavior and learn in an exploratory manner. At the same time, it is a learning model that has the ability to automatically switch between exploration and exploitation modes and the potential to realize higher explorations that reflect what it has learned so far. However, the learning algorithms in CBRL have not been well-established in previous studies and have yet to incorporate recent advances in reinforcement learning. This study introduced Twin Delayed Deep Deterministic Policy Gradients (TD3), which is one of the state-of-the-art deep reinforcement learning algorithms that can treat deterministic and continuous action spaces, to CBRL. The validation results provide several insights. First, TD3 works as a learning algorithm for CBRL in a simple goal-reaching task. Second, CBRL agents with TD3 can autonomously suppress their exploratory behavior as learning progresses and resume exploration when the environment changes. Finally, examining the effect of the agent’s chaoticity on learning shows that extremely strong chaos negatively impacts the flexible switching between exploration and exploitation.
@@ 1 Introduction
Neural networks (NNs) have been studied for many years, partially inspired by findings in neuroscience research [1]. In recent years, the development of deep learning techniques used to successfully train deep NNs has produced remarkable results in various machine learning fields. The recent progress in this field began with outstanding achievements in the study of image recognition with NNs [2, 3, 4]. Recurrent neural networks (RNNs) also performed well in time-series processing, such as speech recognition and natural language processing [5, 6, 7, 8]. Vaswani et al. proposed the transformer, which is an innovative model for natural language processing [9]. The transformer has provided a breakthrough in the capabilities of artificial intelligence (AI) as the underlying technology for Large Language Models (LLMs) [10]. It has also demonstrated high performance in other areas such as image processing and speech recognition [11, 12]. AI capabilities based on deep learning techniques have developed quickly over the past decade, leading to innovations in diverse fields.
@@
In recent years, AI has demonstrated even high-quality creative abilities [13], and the creativity of AI has been attracting much attention [14, 15]. The outputs of LLMs can be made more conservative or more creative by modifying a probability distribution parameter called “temperature” [16]. Historically, J. McCarthy et al. discussed the difference between creative thinking and unimaginative competent thinking in a proposal for the Dartmouth Workshop in 1956 [17]. They raised the importance of randomness in AI and the need to pursue how randomness can be effectively injected into machines, just as the brain does. In reinforcement learning, a learning agent performs exploratory action driven by random numbers to an environment and improves its policy based on feedback from the environment. Reinforcement learning is inspired by psychological findings and the principle of reinforcement, which states that agents learn from their actions and their consequences. Various studies on reinforcement learning have been conducted over long years [18, 19, 20, 21, 22, 23]. In recent years, research on deep reinforcement learning, which incorporates deep learning techniques, has become popular, and this approach has made it possible to learn even more difficult tasks [24, 25, 26, 27, 28]. Deep reinforcement learning can also provide effective learning performance in tasks such as Go and video games, where it is difficult to specify the desired behavior designed by human hands [29, 30, 31]. Randomness enables AI systems to generate creative outputs and perform exploratory learning.
@@
This study focuses on the spontaneous and autonomous exploration of organisms. Organisms can act spontaneously and autonomously in environments with diverse dynamics, and can adapt to the environment based on the experience gained through these behavior [18]. Reinforcement learning agents stochastically explore the environment by introducing random numbers to select actions that deviate from the optimal action determined by the current policy. On the other hand, it remains a fascinating unsolved problem to understand how the biological brain behaves in various ways and realizes exploratory learning. One hypothesis for the essential property of the source of exploration is to utilize fluctuations within the brain. Various studies have shown that fluctuations caused by spontaneous background activity in the neural populations vary their responses [32]. Briggman et al. demonstrated that decision-making of escape behavior in leeches depends on the activity of specific neurons that fluctuate under the influence of neuronal population dynamics [33]. Fox et al. measured activity in the human motor cortex with fMRI while performing a button-pressing task, and they found a correlation between fluctuations in the BOLD signals and behavioral variability [34]. These studies suggest that fluctuations in neural activity influence decision-making and produce a variety of behaviors.
@@
Freeman pointed out the possibility that the brain uses chaotic dynamics for exploratory learning [35]. Skarda and Freeman observed the EEG of the rabbit olfactory system and showed that there are many chaotic attractors in the dynamics of the olfactory bulb that are attracted when the rabbit is exposed to the known olfactory conditioned stimulus [36]. This study has also confirmed that when the rabbit is exposed to an unknown odorant, chaotic dynamics that itinerate between existing attractors emerge, followed by the reorganization of existing attractors and the emergence of new attractors corresponding to new stimuli. Freeman argues that the chaotic dynamics of the brain continually generate new patterns of activity necessary to generate new structures and that this underlies the ability of trial and error problem solving [35]. Aihara et al. discovered the chaotic dynamics in the squid giant axon and constructed chaotic neural network models based [37, 38, 39], and proposed Chaotic Simulated Annealing effective for combinatorial optimization problems [40]. Hoerzer et al. showed that a reservoir network, which fixes the recurrent and input weights, can acquire various functions using an exploratory learning algorithm based on random noise [41]. This study examined how generic neural circuits that resemble a cortical column can acquire and maintain computational functions, including memory attractors, through biologically plausible learning rules rather than supervised learning. The result suggests that stochasticity (trial-to-trial variability) of neuronal response plays a crucial role in the learning process. It has also been shown that the system’s own chaotic dynamics can drive exploration in this learning process [42, 43]. These studies have implications for understanding how the brain achieves exploratory learning and utilizes chaotic dynamics in the process.
Shibata et al. have proposed chaos-based reinforcement learning (CBRL), which exploits the internal chaotic dynamics of the system for exploration [44]. This method uses an RNN as an agent system and its internal chaotic dynamics as a source of exploration components. Shibata et al. hypothesize that the system can acquire transient dynamics with higher functions by developing its dynamics purposive through exploratory learning. Various studies have viewed brain activity as transitive or transient dynamics [45, 46, 47]. CBRL provides insights from a reinforcement learning perspective on studies of the exploratory nature and the transient dynamics of the brain. CBRL agents also have the advantage that their exploration can autonomously subside or resume depending on the situation. It has been suggested that CBRL agents can acquire higher exploration strategies that reflect their learned behavior [48]. CBRL offers a new perspective on both understanding the brain’s exploratory learning mechanisms and novel learning algorithms for AI.
@@
The algorithms used in CBRL require treating deterministic and continuous action to eliminate stochastic exploration and human-designed action selection. Due to the necessity of a reinforcement learning algorithm that can handle deterministic and continuous actions without requiring random noise, Shibata et al. proposed causality trace learning and used it to train CBRL agents [44, 48, 49, 50]. However, this method has limited performance, and the CBRL algorithm is not yet well established. This study introduces the Twin Delayed Deep Deterministic Policy Gradient (TD3) algorithm, which is one of the state-of-the-art deep reinforcement learning algorithms designed for handling deterministic and continuous actions. The ability of the agent is evaluated by learning a simple goal task. We also examine how CBRL agents trained using TD3 respond to environmental changes. Furthermore, we investigate how the behavior of the agents changes depending on the levels of chaoticity in the model’s dynamics and how this affects their learning performance and ability to adapt to environmental changes.
@@
This paper is organized as follows. Section 2 summarizes chaos-based reinforcement learning. Section 3 describes the experimental method. Section 4 presents the results of the experiments. Section 5 summarizes the conclusions of this study and discusses future research directions.
@@
Fig. 1: Chaos-Based Reinforcement Learning (CBRL). (a) Overview of Regular Reinforcement Learning: The agent decides an action, and then stochastic choices or noise based on random numbers affect the action and drive exploration. The action changes the agent’s state through the environment. The agent improves its policy based on experience gained from the interactions. (b) Overview of CBRL: A dynamical system that exhibits chaotic behavior is used as the agent’s system. In CBRL, the agent behaves exploratively due to fluctuations caused by internal dynamics rather than external random numbers.
@@ 2.2 Expectation for CBRL
Shibata et al. hypothesized that the key to creating systems with higher exploratory behavior and human-like intelligence is having a dynamical system with rich and spontaneous activity [44]. The dynamics of brain activity can be considered as a chaotic itinerancy [45] or as a process that transits through a series of saddle points [46]. The hypothesis expects the system to reconstruct its exploratory chaotic dynamics into transient dynamics that is purposeful to maximize rewards.
Since internal dynamics drives exploration, CBRL agents are also expected to be able to optimize exploration itself through learning. Goto et al. demonstrated that CBRL agents can change motor noise-like exploration into more sophisticated exploration that selects routes to avoid obstacles using an obstacle avoidance task [48, 49]. It is expected that CBRL agents can acquire more advanced exploration capabilities that effectively utilize previously acquired behaviors by constructing the agent’s internal state and decision-making processes as transitive dynamics such as chaotic itinerancy.
@@ 2.3 Issue of learning algorithm for CBRL
To ensure freedom and autonomy in learning, CBRL agents have used algorithms that deal with deterministic and continuous action spaces rather than stochastic and discrete ones in which action selections are defined heteronomously. In previous studies, the learning algorithm for CBRL agents has been an Actor-Critic, which can handle deterministic and continuous action spaces [44]. However, a classic Actor-Critic method that trains an actor using a correlation between the external exploration noise and the resulting change in value cannot be employed for CBRL, which does not use random number exploration. Therefore, causality trace learning, which is similar to Hebbian learning, has been proposed and employed as the training method for the actor-network [44, 48, 49, 50]. This method generates the teacher signal for the actor-network by multiplying the TD error by the input values stored in a tracing mechanism that changes the input tracing rate according to changes in the neuron’s output. However, this method has many problems, such as the inability to introduce the backpropagation method and the difficulty of learning complex tasks. Therefore, the learning algorithm for CBRL has not yet been sufficiently well-established.
@@ 3.1 TD3
In the following part of this subsection, we first describe the DDPG and then the improvements introduced in TD3. In the DDPG, the agent model consists of an actor-network $\mu(s|\theta^{\mu})$ and a critic network $Q(s,a|\theta^{Q})$. Here, $\theta^{\mu}$ and $\theta^{Q}$ are the weight value parameters of each network. $\mu(s|\theta^{\mu})$ determines the action output $a$ based on the agent’s state $s$ and $Q(s,a|\theta^{Q})$ estimates the state action value from $s$ and $a$. Target networks $\mu^{\prime}(s|\theta^{\mu^{\prime}})$ and $Q^{\prime}(s,a|\theta^{Q^{\prime}})$ are generated for each network with their weights copied and are used to generate the teacher signals to stabilize learning.
@@
where the discount rate $0\leq\gamma\leq 1$ is a hyperparameter that determines the present value of future rewards. We then update $\theta^{Q}$ to minimize the loss function such that
@@
The actor-network $\mu$ learns with deterministic policy gradient [54] estimated based on the sampled data:
@@
TD3 introduces three methods to the DDPG algorithm: Clipped Double Q-learning, Delayed Policy Updates, and Target Policy Smoothing. Clipped Double Q-learning is a method to suppress overestimaton of the value by preparing two Critic networks $Q_{1}$ and $Q_{2}$ and adopting the smaller output value when generating teacher signals. Target Policy Smoothing is a technique to suppress the overfitting of $Q$ to inaccurate value estimations by adding noise limited between $-C$ and $C$ to the output of $\mu^{\prime}$ during the generation of the teacher signal. With these techniques, the teacher signal for $Q$ is given by
@@ 3.2 Reservoir network
CBRL requires the agent system to have chaotic dynamics. RNNs are dynamical systems and appropriate models for CBRL agents. However, training RNNs with CBRL presents a challenge in balancing the formation of convergent dynamics beneficial for completing the task while maintaining the chaotic dynamics required for exploration. To avoid this problem, we use a reservoir network (RN) that can be trained without modifying the parameters that determine the dynamical properties of the recurrent layer. Maass and Jaeger have proposed RNs as a Liquid State Machine (LSM) and an Echo State Network (ESN), respectively [55, 56]. In this study, we use an ESN composed of rate model neurons. The ESN consists of a recurrent layer called the “reservoir” and an output layer called the “readout.” The connection weights in the reservoir are randomly initialized and fixed, and only the weights from the reservoir to the readout are trained. The reservoir receives the time-series input and generates an output that nonlinearly reflects the spatio-temporal context of the input series. The readout generates the network output by performing a linear combination of the reservoir output and input. The dynamical properties of the ESN can be modified with a single fixed parameter. Therefore, it is easy to tune the chaoticity of the system during learning with chaos-based exploration [43]. Note that, in this study, the ESN is not used to process time series data but rather to generate chaotic dynamics in the agent system.
@@
The structure of the ESN is shown in Fig. 2. The reservoir has $N_{\rm{x}}$ neurons that are recurrently connected with $\bm{W}^{\rm{rec}}\in\mathbb{R}^{N_{\rm{x}}\times N_{\rm{x}}}$ with probability of $p$. The reservoir receives $N_{\rm{i}}$-dimensional inputs through the weight matrix $\bm{W}^{\rm{in}}\in\mathbb{R}^{N_{x}\times N_{i}}$. Then, the reservoir output is computed by the following equation:
@@
where $g$ is a scaling parameter that scales the strength of the recurrent connections and $\bm{u}_{t}\in\mathbb{R}^{N_{\rm{i}}}$ is the input vector. $f(\cdot)=\tanh(\cdot)$ is the activation function applied element-wise. Typically, $\bm{W}^{\rm{rec}}$ is a sparse and fixed weight matrix initialized with a spectral radius of 1 by the following procedure. A random $N_{\rm{x}}\times N_{\rm{x}}$ matrix $\bm{W}$ is generated (from a uniform distribution in this study), and then the elements of $\bm{W}$ are set to 00 with probability of $(1-p)$. The matrix is normalized by its spectral radius $\rho$. Thus, the $\bm{W}^{\rm{rec}}$ is initialized as follows
@@
where $\bm{W}^{\rm{out}}\in\mathbb{R}^{N_{\rm{o}}\times(N_{\rm{x}}+N_{\rm{i}})}$ is the output weight matrix, $N_{\rm{o}}$ is the output dimension, and [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes the concatenation of two column vectors. $\bm{W}^{\rm{out}}$ is often fitted using algorithms such as ridge regression in reservoir computing. This study uses a reservoir network as an actor-network for CBRL agents. Since $\bm{W}^{\rm{out}}$ is trained by a slightly modified version of the TD3 algorithm as described in the folloing subsection, $\bm{W}^{\rm{out}}$ is updated by using gradient descent with the Adam optimizer [53].
@@
Fig. 3: Network Configurations. This study compares three network configurations of agents. (a) A TD3-CBRL agent, which uses a chaotic ESN in the actor-network $\mu$. (b) An agent whose actor-network is an MLP with a hidden layer consisting of the same number of neurons as the reservoir in (a). In this case, the agent learns without the external random noises $\epsilon^{\mathrm{t}}_{a},\epsilon^{\mathrm{t}}_{i}$ as added in the regular TD3. (c) An agent that has the same structure as (b) but trains using external random noise as in the regular TD3.

### v2

@@ Abstract
Chaos-based reinforcement learning (CBRL) is a method in which the agent’s internal chaotic dynamics drives exploration. However, the learning algorithms in CBRL have not been thoroughly developed in previous studies, nor have they incorporated recent advances in reinforcement learning. This study introduced Twin Delayed Deep Deterministic Policy Gradients (TD3), which is one of the state-of-the-art deep reinforcement learning algorithms that can treat deterministic and continuous action spaces, to CBRL. The validation results provide several insights. First, TD3 works as a learning algorithm for CBRL in a simple goal-reaching task. Second, CBRL agents with TD3 can autonomously suppress their exploratory behavior as learning progresses and resume exploration when the environment changes. Finally, examining the effect of the agent’s chaoticity on learning shows that there exists a suitable range of chaos strength in the agent’s model to flexibly switch between exploration and exploitation and adapt to environmental changes.
@@ 1 Introduction
Neural networks have been studied for many years, partially inspired by findings in neuroscience research [1]. In recent years, the development of deep learning techniques used to successfully train deep neural networks has produced remarkable results in various fields, e.g. image recognition [2, 3], speech recognition [4, 5], natural language processing [6, 7], and practical applications including next generation wireless communications, economic forecasting, and other areas [8, 9]. The transformer [10], a groundbreaking model for natural language processing, has provided a breakthrough in the capabilities of artificial intelligence (AI) as the underlying technology for Large Language Models (LLMs) [11]. It has also demonstrated high performance in other areas such as image processing and speech recognition [12, 13]. AI capabilities based on deep learning techniques have developed quickly over the past decade, leading to innovations in diverse fields.
@@
In recent years, AI has demonstrated even high-quality creative abilities [14], and the creativity of AI has been attracting much attention [15, 16]. Historically, J. McCarthy et al. discussed the difference between creative thinking and unimaginative competent thinking in a proposal for the Dartmouth Workshop in 1956 [17]. They raised the importance of randomness in AI and the need to pursue how randomness can be effectively injected into machines, just as the brain does. Reinforcement learning is a machine learning approach using random exploration inspired by behavioral insights, where animals learn from their actions and their consequences. In reinforcement learning, a learning agent performs exploratory action driven by random numbers to an environment and improves its policy based on feedback from the environment. Various studies on reinforcement learning have been conducted over long years [18]. In recent years, research on deep reinforcement learning, which incorporates deep learning techniques, has become popular, and this approach has made it possible to learn even more difficult tasks [19, 20] Deep reinforcement learning can also provide effective learning performance in more complex tasks such as Go [21, 22]. Randomness enables AI systems to perform exploratory and interactively learning without explicit teacher.
@@
Organisms can act spontaneously and autonomously in environments with diverse dynamics, and can adapt to the environment based on the experience gained through these behaviors [18]. Reinforcement learning agents stochastically explore the environment by introducing random numbers to select actions that deviate from the optimal action determined by the current policy. On the other hand, it remains a fascinating unsolved problem to understand how the biological brain behaves in various ways and realizes exploratory learning. One hypothesis for the essential property of the source of exploration is to utilize fluctuations within the brain. Various studies have shown that fluctuations caused by spontaneous background activity in the neural populations vary their responses [23]. Experiments measuring neural activity in the leeches [24] and human motor cortex [25] suggest that fluctuations in neural activity influence decision-making and behavioral variability.
@@
Freeman pointed out the possibility that the brain uses chaotic dynamics for exploratory learning [26]. Skarda and Freeman showed that there are many chaotic attractors in the dynamics of the olfactory bulb that are attracted when the rabbit is exposed to the known olfactory conditioned stimulus [27]. This study also suggests that chaotic dynamics is used for the reorganization of new attractors corresponding to new stimuli. Freeman argues that the chaotic dynamics of the brain continually generate new patterns of activity necessary to generate new structures and that this underlies the ability of trial and error problem solving [26]. Aihara et al. discovered the chaotic dynamics in the squid giant axon and constructed chaotic neural network models based [28, 29, 30], and proposed Chaotic Simulated Annealing effective for combinatorial optimization problems [31]. Hoerzer et al. showed that a reservoir network, which fixes the recurrent and input weights, can acquire various functions using an exploratory learning algorithm based on random noise [32]. Additionally, this study suggests that stochasticity (trial-to-trial variability) of neuronal response plays a crucial role in the learning process. It has also been shown that the system’s own chaotic dynamics can drive exploration in this learning process [33, 34]. These studies have implications for understanding how the brain achieves exploratory learning and utilizes chaotic dynamics in the process.
@@
Shibata et al. have proposed chaos-based reinforcement learning (CBRL), which exploits the internal chaotic dynamics of the system for exploration [35]. This method uses a recurrent neural network as an agent system and its internal chaotic dynamics as a source of exploration components. The algorithms used in CBRL require treating deterministic and continuous action to eliminate stochastic exploration and human-designed action selection. Due to the necessity of a reinforcement learning algorithm that can handle deterministic and continuous actions without requiring random noise, Shibata et al. proposed causality trace learning and used it to train CBRL agents [36, 37]. However, this method has limited performance, and the CBRL algorithm is not yet well established. To address this limitation, this study proposes a novel CBRL framework that introduces the Twin Delayed Deep Deterministic Policy Gradient (TD3) algorithm, which is one of the state-of-the-art deep reinforcement learning algorithms designed for handling deterministic and continuous actions. The ability of the agent is evaluated by learning a simple goal task. We also examine how CBRL agents trained using TD3 respond to environmental changes. Through these experiments, we demonstrate that internal-dynamics-based exploration enables more flexible switching between exploration and exploitation, in contrast to standard deterministic approaches that rely on external random noise. Furthermore, we investigate how the behavior of the agents changes depending on the levels of chaoticity in the model’s dynamics and how the dynamics affects their learning performance and ability to adapt to environmental changes. Finally, we evaluate the learning performance of the proposed method with several environments, including a nonlinear control task and a partially observable task.
@@
This paper is organized as follows. Section 2 summarizes chaos-based reinforcement learning. Section 3 describes the experimental method. Section 4 presents the results of the experiments. Section 5 discusses the experimental results. Section 6 summarizes the conclusions of this study and future research directions.
@@
Fig. 1: Chaos-Based Reinforcement Learning (CBRL). (a) Overview of Regular Reinforcement Learning: The agent decides an action, and then stochastic choices or noise based on random numbers affect the action and drive exploration. The action changes the agent’s state through the environment. The agent improves its policy based on experience gained from the interactions. Since exploration noise is externally provided with random numbers, Regular RL agents cannot improve their exploration through learning. (b) Overview of CBRL: A dynamical system that exhibits chaotic behavior is used as the agent’s system. In CBRL, the agent behaves exploratively due to fluctuations caused by internal dynamics rather than external random numbers. Since internal dynamics drives exploration, CBRL agents have the potential to be able to optimize exploration by purposefully adapting their internal dynamics through learning.
@@ 2.2 Expectation for CBRL
CBRL has been studied with the expectation of clarifying the role of chaotic dynamics in biological brains and realizing intelligent systems based on transient dynamics. Shibata et al. hypothesized that the key to creating systems with higher exploratory behavior and human-like intelligence is having a dynamical system with rich and spontaneous activity [35]. The dynamics of brain activity can be considered as a chaotic itinerancy or as a process that transits through a series of saddle points [40, 41, 42]. The hypothesis expects the system to reconstruct its exploratory chaotic dynamics into transient dynamics that is purposeful to maximize returns.
Since internal dynamics drives exploration, CBRL agents are also expected to be able to optimize exploration itself by reconstructing their internal dynamics into purposeful transitions through learning. Goto et al. demonstrated that CBRL agents can change motor noise-like exploration into more sophisticated exploration that selects routes to avoid obstacles using an obstacle avoidance task [43, 36]. It is expected that CBRL agents can acquire more advanced exploration capabilities that effectively utilize previously acquired behaviors by constructing the agent’s internal state and decision-making processes as transitive dynamics such as chaotic itinerancy.
@@ 2.3 Issue of learning algorithm for CBRL
To ensure freedom and autonomy in learning, CBRL agents have used algorithms that deal with deterministic and continuous action spaces rather than stochastic and discrete ones in which action selections are defined heteronomously. In previous studies, the learning algorithm for CBRL agents has been an Actor-Critic, which can handle deterministic and continuous action spaces [35]. However, a classic Actor-Critic method that trains an actor network using a correlation between the external exploration noise and the resulting change in value cannot be employed for CBRL, which does not use random number exploration. Therefore, causality trace learning, which is similar to Hebbian learning, has been proposed and employed as the training method for the actor network [35, 43, 36, 37]. This method generates the teacher signal for the actor network by multiplying the TD error by the input values stored in a tracing mechanism that changes the input tracing rate according to changes in the neuron’s output. However, this method has many problems, such as the inability to introduce the backpropagation method and the difficulty of learning complex tasks. Therefore, the learning algorithm for CBRL has not yet been sufficiently well-established.
@@ 3.1 TD3
In the following part of this subsection, we first describe the DDPG and then the improvements introduced in TD3. In the DDPG, the agent model consists of an actor network $\mu(s|\theta^{\mu})$ and a critic network $Q(s,a|\theta^{Q})$. Here, $\theta^{\mu}$ and $\theta^{Q}$ are the weight value parameters of each network. $\mu(s|\theta^{\mu})$ determines the action output $a$ based on the agent’s state $s$ and $Q(s,a|\theta^{Q})$ estimates the state action value from $s$ and $a$. Target networks $\mu^{\prime}(s|\theta^{\mu^{\prime}})$ and $Q^{\prime}(s,a|\theta^{Q^{\prime}})$ are generated for each network with their weights copied and are used to generate the teacher signals to stabilize learning.
@@
where the discount factor $0\leq\gamma\leq 1$ is a hyperparameter that determines the present value of future rewards. We then update $\theta^{Q}$ to minimize the loss function such that
@@
The actor network $\mu$ learns with deterministic policy gradient [45] estimated based on the sampled data:
@@
TD3 introduces three methods to the DDPG algorithm: Clipped Double Q-learning, Delayed Policy Updates, and Target Policy Smoothing. Clipped Double Q-learning is a method to suppress overestimation of the value by preparing two critic networks $Q_{1}$ and $Q_{2}$ and adopting the smaller output value when generating teacher signals. Target Policy Smoothing is a technique to suppress the overfitting of $Q$ to inaccurate value estimations by adding noise limited between C-C and $C$ to the output of $\mu^{\prime}$ during the generation of the teacher signal. With these techniques, the teacher signal for $Q$ is given by
@@ 3.2 Reservoir network
CBRL requires the agent system to have chaotic dynamics. Recurrent neural networks are dynamical systems and appropriate models for CBRL agents. However, training recurrent neural networks with CBRL presents a challenge in balancing the formation of convergent dynamics beneficial for completing the task while maintaining the chaotic dynamics required for exploration. To avoid this problem, we use an Echo State Network (ESN) that can be trained without modifying the parameters that determine the dynamical properties of the recurrent layer. The ESN is a type of model known as a Reservoir Network [46, 47]. It consists of a recurrent layer called the “reservoir” and an output layer called the “readout.” The connection weights in the reservoir are randomly initialized and fixed, and only the weights from the reservoir to the readout are trained. The reservoir receives the time-series input and generates an output that nonlinearly reflects the spatio-temporal context of the input series. The readout generates the network output by performing a linear combination of the reservoir state and input. The dynamical properties of the ESN can be modified with a single fixed parameter. Therefore, it is easy to tune the chaoticity of the system during learning with chaos-based exploration [34]. Note that, in this study, the ESN is not used to process time series data but rather to generate chaotic dynamics in the agent system.
@@
The structure of the ESN is shown in Fig. 2. The reservoir has $N_{\rm{x}}$ neurons that are recurrently connected with $\bm{W}^{\rm{rec}}\in\mathbb{R}^{N_{\rm{x}}\times N_{\rm{x}}}$ with probability of $p$. The reservoir receives $N_{\rm{i}}$-dimensional inputs through the weight matrix $\bm{W}^{\rm{in}}\in\mathbb{R}^{N_{x}\times N_{i}}$. Then, the reservoir state is computed by the following equation:
@@
where $g$ is a scaling parameter that scales the strength of the recurrent connections and $\bm{u}_{t}\in\mathbb{R}^{N_{\rm{i}}}$ is the input vector. $f(\cdot)=\tanh(\cdot)$ is the activation function applied element-wise. Typically, $\bm{W}^{\rm{rec}}$ is a sparse and fixed weight matrix initialized with a spectral radius of 1 by the following procedure. A random $N_{\rm{x}}\times N_{\rm{x}}$ matrix $\bm{W}$ is generated (from a uniform distribution in this study), and then the elements of $\bm{W}$ are set to 0 with probability of (1p)(1-p). The matrix is normalized by its spectral radius $\rho$. Thus, the $\bm{W}^{\rm{rec}}$ is initialized as follows:
@@
where $\bm{W}^{\rm{out}}\in\mathbb{R}^{N_{\rm{o}}\times(N_{\rm{x}}+N_{\rm{i}})}$ is the output weight matrix, $N_{\rm{o}}$ is the output dimension, and $\cdot$$\cdot$] denotes the concatenation of two column vectors. $\bm{W}^{\rm{out}}$ is often fitted using algorithms such as ridge regression in reservoir computing. This study uses a reservoir network as an actor network for CBRL agents. Since $\bm{W}^{\rm{out}}$ is trained by a slightly modified version of the TD3 algorithm as described in the following subsection, $\bm{W}^{\rm{out}}$ is updated by using gradient descent with the Adam optimizer [44].
@@
Fig. 3: Network Configurations. This study compares three network configurations of agents. (a) A TD3-CBRL agent, which uses a chaotic ESN in the actor network $\mu$. (b) An agent whose actor network is an MLP with a hidden layer consisting of the same number of neurons as the reservoir in (a). In this case, the agent learns without the external random noises $\epsilon^{\mathrm{t}}_{a},\epsilon^{\mathrm{t}}_{i}$ as added in the regular TD3. (c) An agent that has the same structure as (b) but explores using external random noise as in the regular TD3.